3. A composite beam consists of a concrete slab that has an effective width be 72-in and a thickness t 7-in. Compressive strength of concrete is 4 ksi. The slabis connected to a A992 steel beam (W36x160) with an adequate number of steel studs to produce a fully composite beam. Determine the nominal moment strength of this beam.

Please refer AISC 15th edition
Answer #1

effective width of beam = 72 in

thickness of slab= 7 in

grade of concrete = 4 ksi

Beam section = W36x160

Area of beam = 47 in2

depth of beam = 36 in

beam yield strength = 50 ksi

maximum tension that can be taken by beam = 50*47=2350 kips

maximum compression that can be taken by concrete = 0.85*4*72*7=1713.6 kips

Since compression is greater than tension,for full composite action there will be some part of steel in tension

Let area of steel in compression be Ac and that in tension be At

1713.6+50*Ac = 50*At = 50*(47-Ac)=2350-50Ac

Ac = 6.364 in2

At = 47-6.364=40.636 in2

width of flange of W36x160 = 12 in

thickness of flange in compression = 6.364/12=0.53 in

compression force in steel = 50*6.364=318.2 kips

Total compression force = 1713.6+318.2=2031.8 kips

approximate location of resultant tension force from top of steel = 36/2=18 in

lever arm length between compression in concrete and tension in steel = (7/2)+18=21.5 in

lever arm length between compression in steel and tension in steel = 18-0.53/2=17.735 in

nominal moment strength = 1713.6*21.5/12 + 318.2*17.735/12 = 3540 kip-ft